If an object of mass 4kg goes round a circle of a radius 0.5m in 3.142, what is the force towards the centre(π=3.142)


1
Expert's answer
2020-08-19T13:55:05-0400

Th force towards the centre is the centripetal force

"F=\frac{mV^{2}}{r}"

The distance round a circle is the circumference of a circle "2\pi"

Time taken to go round circle "=3.142s"

Velocity of object

"=distance/time=V =\frac{2\pi}{3.142}"


"=\frac{2\times3.142}{3.142}=2ms^{-1}"


"F=\frac{mV^{2}}{r}"


"=\frac{4\times2^{2}}{0.5}"


"=32N"